Please Help with this one I have been trying to figure this out for a while.
A business invests $10,000 in a savings account for two years. At the beginning of the second year, an additional $3500 is invested. At the end of the second year, the account balance is $15,569.75. What was the annual interest rate?

Answer by scott8148(6628)

let r="annual interest rate"

after one year A=10000(1+r) __ A=10000+10000r

3500 is added __ after second year A=(13500+10000r)(1+r) __ 15569.75=13500+10000r+13500r+10000r^2

subtracting 15569.75 __ 10000r^2+23500r-2069.75=0

use the quadratic formula to find r __ r=(-23500ħsqrt[23500^2-(4*10000*(-2069.75))])/[2*10000]

r=8.5%